# Towards Open-World Human Action Segmentation Using Graph Convolutional Networks

### v1

@@ Abstract
Human-object interaction segmentation is a fundamental task of daily activity understanding, which plays a crucial role in applications such as assistive robotics, healthcare, and autonomous systems. Most existing learning-based methods excel in closed-world action segmentation, they struggle to generalize to open-world scenarios where novel actions emerge. Collecting exhaustive action categories for training is impractical due to the dynamic diversity of human activities, necessitating models that detect and segment out-of-distribution actions without manual annotation. To address this issue, we formally define the open-world action segmentation problem and propose a structured framework for detecting and segmenting unseen actions. Our framework introduces three key innovations: 1) an Enhanced Pyramid Graph Convolutional Network (EPGCN) with a novel decoder module for robust spatiotemporal feature upsampling. 2) Mixup-based training to synthesize out-of-distribution data, eliminating reliance on manual annotations. 3) A novel Temporal Clustering loss that groups in-distribution actions while distancing out-of-distribution samples.
We evaluate our framework on two challenging human-object interaction recognition datasets: Bimanual Actions and 2 Hands and Object (H2O) datasets. Experimental results demonstrate significant improvements over state-of-the-art action segmentation models across multiple open-set evaluation metrics, achieving $16.9\%$ and $34.6\%$ relative gains in open-set segmentation (F1@50) and out-of-distribution detection performances (AUROC), respectively. Additionally, we conduct an in-depth ablation study to assess the impact of each proposed component, identifying the optimal framework configuration for open-world action segmentation.
@@ I INTRODUCTION
Recently, Graph Convolutional Networks (GCNs) have presented promising results of action segmentation, particularly through skeleton-based representations that offer robustness to occlusions and computational efficiency [1, 2]. The Pyramid Graph Convolutional Network (PGCN) [1] improves frame-wise action segmentation through multi-scale feature fusion. However, existing models operate in closed-world settings, where training and testing datasets share the same action categories. This assumption does not hold in real-world scenarios, where models frequently encounter unseen actions, leading to poor generalization in open-world settings.
Current approaches to open-world recognition face critical limitations in reflecting true generalization. Existing methods like the Nearest Non-Outlier (NNO) algorithm [3] depend on human-annotated unknown samples during fine-tuning, while ActionCLIP [4] leverages CLIP to extend recognition to unseen actions, which is a large pretrained visual-language model with inherent semantic knowledge of unknown classes. Although effective, both strategies introduce biases: NNO assumes unrealistic access to labeled unknowns, and ActionCLIP leverages external knowledge from pretrained models. These create an unfair advantage that diverges from real-world open-world constraints. Recent advances, such as the Uncertainty-Quantified Temporal Fusion Graph Convolutional Network (UQ-TFGCN) [5], address Out-of-Distribution (OOD) detection by preserving physical distance in the feature space but overlook inter-class discriminability among OOD actions.
To address these issues, we redefine open-world action segmentation as a generalization task where models must recognize and segment unseen actions using only knowledge from training on closed-world classes, eliminating dependencies on external annotations or pretrained language models, as shown in Fig 1. We discard the traditional softmax-based classification output for unknown actions classification, which inherently biases predictions toward known classes due to its closed-world confidence calibration. Instead, we explore the relations between feature space and Out-of-Distribution classification, propose a novel Temporal Efficient Upsampling decoder that enables explicit modeling of fine-grained relationships between different actions, and leverages K-means clustering on feature embeddings to categorize OOD actions. In doing so, we propose a Temporal Clustering Loss to enforce tighter feature grouping along the temporal dimension, improving the model’s ability to distinguish In-Distribution from OOD samples. Furthermore, we incorporate Mixup-based augmentation to expose the model to diverse scenarios during training. These innovations ensure more effective OOD detection and classification, enabling robust action segmentation in open-world scenarios.
@@
  • We conduct extensive experiments on two challenging HOI datasets—Bimanual Actions (Bimacs) [6] and H2O [7] and demonstrate that our approach consistently outperforms existing baseline. Additionally, we perform detailed ablation studies to assess the effectiveness of our framework and evaluate alternative design choices.
@@ II-A Graph Convolutional Network
Recently, Graph Convolutional Networks (GCNs) have emerged as a powerful deep learning framework for learning from graph-structured data. Graphs represent non-Euclidean data structures, and conventional neural networks such as Multilayer Perceptrons (MLPs), Convolutional Neural Networks (CNNs), or Recurrent Neural Networks (RNNs) are not inherently designed to effectively learn from graph data, as they are primarily tailored for Euclidean data (e.g., images, text, RGB-D videos). There are two main types of GCNs commonly used: 1) Spectral GCNs, which operate in the spectral domain. Bruna et al. [8] introduced spectral networks, generalizing CNNs to graphs by leveraging the graph Laplacian spectrum. 2) Spatial GCNs, which operate directly on the graph domain (nodes and edges). For instance, Hamilton et al. [9] introduced fixed aggregation functions to summarize neighborhood features, while Veličković et al. [10] proposed Graph Attention Networks (GATs), extending GCNs by incorporating attention mechanisms. In the context of skeleton-based action recognition, including our work, the latter approach is predominantly adopted.
Yan et al. [2] introduced Spatial-Temporal Graph Convolutional Networks (ST-GCN) for skeleton-based action recognition by leveraging spatio-temporal graphs. Shi et al. [11] extended this with Two-Stream Adaptive Graph Convolutional Networks (2s-AGCN), which adaptively learn joint and bone features. More recently, Myung et al. [12] introduced Deformable Graph Convolutional Networks, which dynamically learn the most informative joint features in both spatial and temporal domains.
@@ II-B Action Segmentation
Temporal action segmentation (TAS) divides an untrimmed video into segments, each assigned an action label, akin to semantic segmentation but in the temporal domain. It enables automatic action recognition by identifying action onset, progression, and conclusion. Approaches in TAS generally follow three architectural designs: (i) Encoder-decoder architectures, such as the Temporal Convolutional Network (TCN) by Lea et al. [13], which leverages temporal convolutions to model long-range dependencies efficiently. (ii) Multistage architectures, exemplified by the Multi-Stage Temporal Convolutional Network (MS-TCN) by Farha et al. [14], which refines predictions iteratively through stacked single-stage TCNs. Filtjens et al. [15] extended this with MS-GCN, integrating spatial-temporal graph convolutions for skeleton-based inputs. (iii) Transformer-based architectures, such as ASFormer by Yi et al. [16], which employs self-attention in the encoder and cross-attention in the decoder to capture complex temporal dependencies. This work follows the encoder-decoder architecture for our model design.
@@ II-C Human-object interaction (HOI) recognition
Video-based human-object interaction (HOI) recognition analyzes the temporal structure of untrimmed videos to identify sub-activities and object affordances. Traditional methods, such as Conditional Random Fields (CRFs), have been largely replaced by deep learning approaches, including CNNs, RNNs, and 3D CNNs, due to their superior ability to model complex relationships. More recently, Graph Convolutional Networks (GCNs) have gained traction for capturing spatial and temporal dependencies in HOI tasks. Morais et al. [17] introduced the Asynchronous-Sparse Interaction Graph Network (ASSIGN), which models HOI as a spatio-temporal graph, enabling asynchronous entity updates for improved segmentation but suffering from RNN-related short-term memory limitations. Qian et al. [18] proposed the Two-level Geometric feature informed Graph Convolutional Network (2G-GCN), which fuses geometric skeleton-based representations with RGB video features to mitigate occlusion issues. Their fusion-level network integrates an attention mechanism to enhance interaction modeling, leveraging ASSIGN as the backbone for HOI recognition.
Figure 2: Architecture of the Efficient Pyramid Graph Convolutional Network (EPGCN) with the Temporal Efficient Upsampling (TEU) Module. The TEU decoder processes multi-scale encoder graph features ($G4$, $G7$, $G10$) through dual pathways: (1) the Downsampling Path ($I_{d}\rightarrow A,B\rightarrow C$); (2) the Upsampling Path ($I_{u}\rightarrow D\rightarrow\tilde{D}$). The fused output is processed by a Temporal Pyramid Pooling (TPP) [1] module ($C,\tilde{D},D,\rightarrow E\rightarrow F\rightarrow I$). The produced feature maps $F$ and $I$ are collected for intra/inter-clustering loss optimization and OOD classification, respectively.
@@ II-D Generalized Out-Of-Distribution(OOD) Framework
Out-of-Distribution (OOD) research is critical in deep learning, as models frequently encounter unseen test data that differ from their training distribution, a phenomenon known as distributional shift. This shift is categorized into covariate shift, where ID and OOD samples originate from different domains, and semantic shift, where both share the same domain but belong to different semantic classes. To address these challenges, the generalized OOD framework includes several sub-tasks. (i) OOD detection determines whether a sample belongs to an unknown distribution, with methods such as thresholding softmax probabilities [19], while Liang et al. [20] improved performance by calibrating softmax outputs and applying temperature scaling. (ii) Open-set recognition (OSR) extends OOD detection by also classifying known samples. For instance, Bao et al. [21] proposed Deep Evidential Action Recognition (DEAR), leveraging an Evidential Neural Network (ENN) to predict a Dirichlet distribution over class probabilities. (iii) Generalized zero-shot learning (GZSL) aims to recognize both seen and unseen classes, with approaches such as conditional Wasserstein GANs (WGAN) [22] for feature synthesis, although these require prior knowledge of unknown features. (iv) Open-world recognition (OWR), introduced by Bendale et al. [3], requires models to detect and incrementally learn new classes using human annotation by employing the Nearest Non-Outlier (NNO) algorithm.
Our work follows the OWR paradigm but eliminates the incremental learning stage that relies on human-labeled novel data, which is particularly infeasible for segmentation tasks requiring fine-grained frame-wise labeling. Moreover, we adopt the thresholding approach of Hendrycks et al. [19], using output logits to distinguish OOD from ID samples.
@@ III Methodologies
III-A Open-World Action Segmentation
Problem definition and notation: formally, let the set of labeled action classes observed during training be denoted as $C_{\text{known}}$, meaning that all training samples belong to one of these known classes, expressed as $x\in C_{\text{known}}$. In a conventional Closed-World Action Segmentation (CWAS) setting, the test set consists only of label-known actions, meaning that the set of all test samples remains within the same distribution as the training set, i.e., $C_{\text{all}}=C_{\text{known}}$. However, in the Open-World Action Segmentation (OWAS) setting, an additional set of novel action classes, $C_{\text{novel}}$, exists at inference time. These novel classes are completely disjoint from the known training classes, satisfying $C_{\text{novel}}\cap C_{\text{known}}=\emptyset$. As a result, the set of all test samples in OWAS consists of both known and novel action classes, i.e., $C_{\text{all}}=C_{\text{known}}\cup C_{\text{novel}}$.
Definition: the workflow of the Open-World Action Segmentation problem $[F,x,\phi,\nu,\kappa]$ is given by:
@@ III-A Open-World Action Segmentation
    Clustering and label assignment: For novel samples $\{x|\nu(\phi(x))=0\}$, a clustering function $\kappa$ groups them into $M$ distinct pseudo-classes. These are mapped to incrementally indexed new classes $C_{novel}=\{C_{known}+1,C_{known}+2,\dots\}$ via Hungarian algorithm.
@@ III-B Enhanced Pyramid Graph Convolutional Network
We introduce an Enhanced Pyramid Graph Convolutional Network (EPGCN) for motion feature extraction and open-world action recognition. As illustrated in Fig 2, EPGCN extends the Pyramid Graph Convolutional Network (PGCN) [1] baseline by integrating a Temporal Efficient Upsampling (TEU) decoder, designed to preserve discriminative spatiotemporal features critical for open-world generalization.
The encoder extracts hierarchical motion representations at three resolutions ($G4$, $G7$, $G10$), corresponding to temporal scales of $T$, $T/2$, $T/4$ for an input sequence of length $T$. The TEU decoder processes these multi-scale features through two parallel pathways: downsampling path and umsampling path. The downsampling path aggregates global context by progressively reducing temporal resolution, enhancing inter-class discriminability for robust novelty detection. The umsampling path recovers fine-grained motion details through learned temporal interpolation, preserving intra-class structural consistency for precise segmentation.
@@
Umsampling Path: The high-level ($G^{10}$) and mid-level ($G^{7}$) features are temporally upsampled to align with the low-level feature map ($G^{4}$). These are concatenated as: $I_{u}=[G_{4},G_{7}^{\uparrow},G_{10}^{\uparrow}]\in\mathbb{R}^{N\times C\times T\times V}$. A $1\times 1$ convolution is applied to $I_{u}$ to produce feature map $D$. The feature map $B$ from the downsampling path undergoes temporal average pooling, is replicated across the temporal axis, and fused with $D$ via element-wise addition:
@@
The output from TEU is forwarded to a Temporal Pyramid Pooling (TPP) [1] module to capture global context efficiently. It applies temporal average pooling over hierarchically divided time segments, reducing complexity while preserving key temporal patterns. The final feature maps are collected by the K-means algorithm for OOD classification.

### v2

@@ Abstract
Current methods for human-object interaction segmentation excel in closed-world settings but struggle to generalize to open-world scenarios where novel actions emerge. Since collecting exhaustive training data for all possible dynamic human activities is impractical, a model capable of detecting and segmenting novel, out-of-distribution (OOD) actions without manual annotation is needed. To address this, we formally define the open-world action segmentation problem and propose a novel framework featuring three key components: 1) an Enhanced Pyramid Graph Convolutional Network with a new decoder for robust spatiotemporal upsampling, 2) hybrid-based training synthesizing OOD data to eliminate reliance on manual labels, and 3) a temporal clustering loss that groups in-distribution actions while distancing OOD samples.
We evaluate our framework on two challenging human-object interaction recognition datasets: Bimanual Actions and Two Hands and Object datasets. Experimental results demonstrate significant improvements over state-of-the-art action segmentation models across multiple open-set evaluation metrics, achieving $16.9\%$ and $34.6\%$ relative gains in open-set segmentation (F1@50) and out-of-distribution detection performances (AUROC), respectively. Additionally, we conduct an in-depth ablation study to assess the impact of each proposed component, identifying the optimal framework configuration for open-world action segmentation.
@@ I INTRODUCTION
Recently, Graph Convolutional Networks (GCNs) have presented promising results of action segmentation, particularly through skeleton-based representations that offer robustness to occlusions and computational efficiency [xing2022understanding, Yan2018SpatialTG]. The Pyramid Graph Convolutional Network [xing2022understanding] improves frame-wise action segmentation through multi-scale feature fusion. However, existing models operate in closed-world settings, where training and testing datasets share the same action categories. This assumption does not hold in real-world scenarios, where models frequently encounter novel actions, leading to poor generalization in open-world settings.
Current approaches to open-world recognition face critical limitations in reflecting true generalization. Existing methods like the Nearest Non-Outlier (NNO) algorithm [7298799] depend on human-annotated unknown samples during fine-tuning, while ActionCLIP [Wang2021ActionCLIPAN] leverages CLIP to extend recognition to novel actions, which is a large pretrained visual-language model with inherent semantic knowledge of unknown classes. Although effective, both strategies introduce biases: NNO assumes unrealistic access to labeled unknowns, and ActionCLIP leverages external knowledge from pretrained models. These create an unfair advantage that diverges from real-world open-world constraints. Recent advances, such as the Uncertainty-Quantified Temporal Fusion Graph Convolutional Network [xing2024understanding], address Out-of-Distribution (OOD) detection by preserving physical distance in the feature space but overlook inter-class discriminability among OOD actions.
To address these issues, we redefine open-world action segmentation as a generalization task where models must recognize and segment novel actions using only knowledge from training on closed-world classes, eliminating dependencies on external annotations or pretrained language models, as shown in Fig 1. We discard the traditional softmax-based classification output for unknown actions classification, which inherently biases predictions toward known classes due to its closed-world confidence calibration. Instead, we explore the relations between feature space and Out-of-Distribution classification, propose a novel Temporal Efficient Upsampling decoder that enables explicit modeling of fine-grained relationships between different actions, and leverages K-means clustering on feature embeddings to categorize OOD actions. In doing so, we propose a Temporal Clustering Loss to enforce tighter feature grouping along the temporal dimension, improving the model’s ability to distinguish In-Distribution from OOD samples. Furthermore, we incorporate Mixup-based augmentation to expose the model to diverse scenarios during training. These innovations ensure more effective OOD detection and classification, enabling robust action segmentation in open-world scenarios.
@@
  • We conduct extensive experiments on two challenging HOI datasets—Bimanual Actions (Bimacs) [Dreher2019LearningOR] and H2O [Kwon_2021_ICCV] and demonstrate that our approach consistently outperforms existing baseline. Additionally, we perform detailed ablation studies to assess the effectiveness of our framework and evaluate alternative design choices.
@@ II-A Graph Convolutional Network
Recently, Graph Convolutional Networks (GCNs) have emerged as a powerful deep learning framework for learning from graph-structured data. Graphs represent non-Euclidean data structures, and conventional neural networks (CNNs) such as Multilayer Perceptrons, Convolutional Neural Networks, or Recurrent Neural Networks are not inherently designed to effectively learn from graph data, as they are primarily tailored for Euclidean data (e.g., images, text, RGB-D videos). There are two main types of GCNs commonly used: 1) Spectral GCNs, which operate in the spectral domain. Bruna et al. [Bruna2013SpectralNA] introduced spectral networks, generalizing CNNs to graphs by leveraging the graph Laplacian spectrum. 2) Spatial GCNs, which operate directly on the graph domain (nodes and edges). For instance, Hamilton et al. [Hamilton2017InductiveRL] introduced fixed aggregation functions to summarize neighborhood features, while Veličković et al. [velickovic2018graph] proposed Graph Attention Networks (GATs), extending GCNs by incorporating attention mechanisms. In the context of skeleton-based action recognition, including our work, the latter approach is predominantly adopted.
Yan et al. [Yan2018SpatialTG] introduced Spatial-Temporal Graph Convolutional Networks (ST-GCN) for skeleton-based action recognition by leveraging spatio-temporal graphs. Shi et al. [2sagcn2019cvpr] extended this with Two-Stream Adaptive Graph Convolutional Networks (2s-AGCN), which adaptively learn joint and bone features. More recently, Myung et al. [10478824] introduced Deformable Graph Convolutional Networks, which dynamically learn the most informative joint features in both spatial and temporal domains.
@@ II-B Action Segmentation
Temporal action segmentation (TAS) divides an untrimmed video into segments, each assigned an action label, akin to semantic segmentation but in the temporal domain. It enables automatic action recognition by identifying action onset, progression, and conclusion. Approaches in TAS generally follow three architectural designs: (i) Encoder-decoder architectures, such as the Temporal Convolutional Network (TCN) by Lea et al. [Lea2016TemporalCN], which leverages temporal convolutions to model long-range dependencies efficiently. (ii) Multistage architectures, exemplified by the Multi-Stage Temporal Convolutional Network (MS-TCN) by Farha et al. [AbuFarha2019MSTCNMT], which refines predictions iteratively through stacked single-stage TCNs. Filtjens et al. [9998567] extended this with MS-GCN, integrating spatial-temporal graph convolutions for skeleton-based inputs. (iii) Transformer-based architectures, such as ASFormer by Yi et al. [chinayi_ASformer], which employs self-attention in the encoder and cross-attention in the decoder to capture complex temporal dependencies. This work follows the encoder-decoder architecture for our model design.
@@ II-C Human-object interaction (HOI) recognition
Video-based human-object interaction (HOI) recognition analyzes the temporal structure of untrimmed videos to identify sub-activities and object affordances [xing2021robust]. Traditional methods, such as Conditional Random Fields, have been largely replaced by deep learning approaches, including CNNs, RNNs, and 3D CNNs, due to their superior ability to model complex relationships. More recently, Graph Convolutional Networks have gained traction for capturing spatial and temporal dependencies in HOI tasks. Morais et al. [9577351] introduced the Asynchronous-Sparse Interaction Graph Network (ASSIGN), which models HOI as a spatio-temporal graph, enabling asynchronous entity updates for improved segmentation but suffering from RNN-related short-term memory limitations. Qian et al. [Qiao2022GeometricFI] proposed the Two-level Geometric feature informed Graph Convolutional Network (2G-GCN), which fuses geometric skeleton-based representations with RGB video features to mitigate occlusion issues. Their fusion-level network integrates an attention mechanism to enhance interaction modeling, leveraging ASSIGN as the backbone for HOI recognition.
Figure 2: Architecture of the Efficient Pyramid Graph Convolutional Network (EPGCN) with the Temporal Efficient Upsampling (TEU) Module. The TEU decoder processes multi-scale encoder graph features ($G4$, $G7$, $G10$) through dual pathways: (1) the Downsampling Path ($I_{d}\rightarrow A,B\rightarrow C$); (2) the Upsampling Path ($I_{u}\rightarrow D\rightarrow\tilde{D}$). The fused output is processed by a Temporal Pyramid Pooling (TPP) [xing2022understanding] module ($C,\tilde{D},D,\rightarrow E\rightarrow F\rightarrow I$). The produced feature maps $F$ and II are collected for intra/inter-clustering loss optimization and OOD classification, respectively.
@@ II-D Generalized Out-Of-Distribution(OOD) Framework
Out-of-Distribution (OOD) research is critical in deep learning, as models frequently encounter unseen test data that differ from their training distribution, a phenomenon known as distributional shift. This shift is categorized into covariate shift, where ID and OOD samples originate from different domains, and semantic shift, where both share the same domain but belong to different semantic classes. To address these challenges, the generalized OOD framework includes several sub-tasks. (i) OOD detection determines whether a sample belongs to an unknown distribution, with methods such as thresholding softmax probabilities [Hendrycks2016ABF], while Liang et al. [Liang2017EnhancingTR] improved performance by calibrating softmax outputs and applying temperature scaling. (ii) Open-set recognition extends OOD detection by also classifying known samples. For instance, Bao et al. [BaoICCV2021DEAR] proposed Deep Evidential Action Recognition, leveraging an Evidential Neural Network to predict a Dirichlet distribution over class probabilities. (iii) Generalized zero-shot learning aims to recognize both seen and unseen classes, with approaches such as conditional Wasserstein GANs [Mandal_2019_CVPR] for feature synthesis, although these require prior knowledge of unknown features. (iv) Open-world recognition (OWR), introduced by Bendale et al. [7298799], requires models to detect and incrementally learn new classes using human annotation by employing the Nearest Non-Outlier algorithm.
Our work follows the OWR paradigm but eliminates the incremental learning stage that relies on human-labeled novel data, which is particularly infeasible for segmentation tasks requiring fine-grained frame-wise labeling. Moreover, we adopt the thresholding approach of Hendrycks et al. [Hendrycks2016ABF], using output logits to distinguish OOD from ID samples.
@@ III Methodologies
III-A Open-World Action Segmentation Workflow
Problem definition and notation: formally, let $C_{\text{known}}$ denote the set of labeled action classes during training, where all training samples satisfy $x\in C_{\text{known}}$. In the Closed-World Action Segmentation (CWAS) setting, test samples consist only of label-known actions, where all test samples remain within the same distribution as the training set. Conversely, in the Open-World Action Segmentation (OWAS) setting, test samples include actions from novel classes $C_{\text{novel}}$ that are disjoint from the known training classes $(C_{\text{novel}}\cap C_{\text{known}}=\emptyset)$.
The workflow of the Open-World Action Segmentation problem $x,\phi,F,\nu,\kappa]$ is given by:
@@ III-A Open-World Action Segmentation Workflow
    Clustering and label assignment: For novel samples $\{x|\nu(\phi(x))=0\}$, a clustering function $\kappa$ groups them into distinct pseudo-classes $M$. These are mapped to incrementally indexed new classes $C_{novel}=\{C_{known}+1,C_{known}+2,\dots C_{known}+M\}$ via Hungarian algorithm.
@@ III-B Enhanced Pyramid Graph Convolutional Network
To realize the OWAS workflow, we introduce an Enhanced Pyramid Graph Convolutional Network (EPGCN). As illustrated in Fig 2, EPGCN implements an attention based graph convolutional encoder [xing2022understanding], introduces a Temporal Efficient Upsampling (TEU) decoder combining with Temporal Pyramid Pooling (TPP). The encoder-decoder structure form the mapping function $\phi$. The known actions are predicted by a convolutional layer ($F$), while the novel actions are clustered by a K-means algorithm ($\kappa$).
The encoder extracts hierarchical motion representations at three resolutions ($G4$, $G7$, $G10$), corresponding to temporal scales of $T$, $T/2$, $T/4$ for an input sequence of length $T$. The TEU decoder processes these multi-scale features through two parallel pathways: downsampling path and upsampling path. The downsampling path aggregates global context by progressively reducing temporal resolution, enhancing inter-class discriminability for robust novelty detection. The upsampling path recovers fine-grained motion details through learned temporal interpolation, preserving intra-class structural consistency for precise segmentation.
@@
Upsampling Path: The high-level ($G^{10}$) and mid-level ($G^{7}$) features are temporally upsampled to align with the low-level feature map ($G^{4}$). These are concatenated as: $I_{u}=[G_{4},G_{7}^{\uparrow},G_{10}^{\uparrow}]\in\mathbb{R}^{N\times C\times T\times V}$. A $1\times 1$ convolution is applied to $I_{u}$ to produce feature map $D$. The feature map $B$ from the downsampling path undergoes temporal average pooling, is replicated across the temporal axis, and fused with $D$ via element-wise addition:
@@
The output from TEU is forwarded to a Temporal Pyramid Pooling (TPP) [xing2022understanding] module to capture global context efficiently. It applies temporal average pooling over hierarchically divided time segments, reducing complexity while preserving key temporal patterns. The final feature maps are collected by the K-means algorithm for OOD classification.